Through(3,4) perpendicular to y =-3x+2

Dmitri needs 5 liters of a 43% solution of sulfuric acid for a research project in molecular biology. He has two supplies of sulfuric acid solution: one is an unlimited supply of the 40% solution and the other an unlimited supply of the 45% solution. How many liters of each solution should Dmitri use?

Answer by ewatrrr(24785)


Hi
Through(3,4) perpendicular to y =-3x+2
y = (1/3)x + b Slopes of perpendicular lines are negative reciprocals
(3,4)
4 = 1 + b substituting the ordered pair to solve for b
3 = b
y = (1/3)x + 3
2nd question
Let x represent the 45% solution, then (5-x) would represent the 40% solution
.40(5-x) + .45x = .43*5L
solving for x
.05x = .43*5 - .40*5L
.05x = .03*5 = .15L
x = 3 L of 45% solution and (5-3) or 2L of 40% solution
checking our answer
.40*2 + .45*3 = .08 + 1.35 = 2.15